ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 22, 46 Kosyak, Alexandr, Institute of Mathematics, Kiev, Ukraine; Zekri, Richard, Université de la Mediterranée, Marseille, France On von Neumann algebras, generated by regular representations of infinite-dimensional groups. For infinite-dimensional groups we prove that the commutant of the analog of the right regular representation is generated by the analog of the left regular representation, when the measure, that defines corresponding representation, is quasi-invariant with respect to the inverse mapping. In this case the von Neumann algebra, generated by the analog of the regular representation is a factor for some groups [1]. Reference: [1] A. V. Kosyak and R. Zekri, Type of von Neumann algebras, generated by regular representations of infinite-dimensional groups, Vestnik Tambov Univ. ,(1998), vol. 3, issue 1, 47-48.